OpenStudy (anonymous):

Mary is solving the equation 2^n – 3 = 83. Her first three steps are shown: Step 1: 2n – 3 = 83 Step 2: 2n = 86 Step 3: ln 2n = ln 86 Which shows Step 4? ln 2 = n ln 86 n + ln 2 = ln 86 n ln 2 = ln 86 n = ln 86 – ln 2

Nnesha (nnesha):

\[\huge\rm 2^{n}-3=86\] is this ur question ? n is exponent right ?

Nnesha (nnesha):

power rule \[\large\rm log_b x^y = y \log_b x\]

OpenStudy (anonymous):

yes thank you

Nnesha (nnesha):

\[\huge\rm ln 2^n =\ln 86\] next u should apply the power rule